OpenStudy (anonymous):

What is the equation of this circle? Answer x2 + (y – 1)2 = 4 x2 + (y + 1)2 = 4 x2 + (y – 1)2 = 2 x2 + (y + 1)2 = 2

OpenStudy (anonymous):

What are the given pieces of information?

OpenStudy (anonymous):

A circle is shown on the coordinate plane below. Circle that contains points 2, 1; 0, 3; negative 2, 1; 0, negative 1.

OpenStudy (anonymous):

Because is is just x^2, the x coordinate of the center of the circle must be 0. Which point(s) have that?

OpenStudy (anonymous):

0, 3 and 0, -1 It can't be 0,3 because then the equation of that circle will be of the form x^2 + (y-3)^2 = r^2. So, it has to be 0, -1.

OpenStudy (anonymous):

thank you once again u should be a mod XD

OpenStudy (anonymous):

I don't need the moderator power. :)

OpenStudy (mertsj):

If (0,3) is on the circle and (0,-1) is on the circle, the center of the circle must be half way between. So the center is (0,1) and the equation of the circle is x^2+(y-1)^2 = 4

OpenStudy (anonymous):

I am confused. Were those the choices for centers of circles are "points ON" the circle?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

I do not see how this statement is true: If (0,3) is on the circle and (0,-1) is on the circle, the center of the circle must be half way between.

OpenStudy (anonymous):

yes i see how its true r u talking to me or Mertsj

OpenStudy (anonymous):